# Asymmetric thin-shell wormholes in the Kalb-Ramond background: Observational characteristics and extra photon rings

### v2

@@ II Null Geodesics and Effective Potential of the ATSW in the Kalb-Ramond Field
With Eqs. (11) (12) (13), we can plot the photon trajectories for different values of the charge $Q$ and the Lorentz-violating parameter $l$, as shown in Fig. 3  and Fig. 4 . Taking $Q=0.1$ and $l=0.01$ as an example, Fig.  3 (a), (b), and (c) illustrate three typical types of photon motion. For the case where $Zb_{c_{2}}<b_{1}<b_{c_{1}}$, photons originating from infinity in spacetime $M_{1}$ exhibit an increasing number of trajectory loops in spacetime $M_{2}$ as the impact parameter $b_{i}$ decreases. By comparing the corresponding columns in Fig. 3  and Fig. 4 , we observe that $b_{c_{1}}$ decreases with an increase in either $Q$ or $l$, while $Zb_{c_{2}}$ shows the opposite trend. Furthermore, it is evident that the photon trajectories are more significantly influenced by the Lorentz-violating parameter $l$.
@@ III.1 The trajectory of a photon
First, in a stationary coordinate system with the BH fixed in the equatorial plane, for a static observer located at infinite distance in the direction of the North Pole, the light source comes from infinite distance in the direction of the South Pole. By categorizing the total number n of times photon trajectories cross the equatorial plane, Gralla et al. studied the optical appearance of a BH surrounded by a thin disk Luo:2023wru
@@
Here, $n_{2}$ and $n_{3}$ serve as additional orbit functions for the ATSW, corresponding to additional photons. This situation is analogous to that of a BH. To better illustrate the relationship between the number of photon trajectories and the impact parameter, we plotted Fig. 5  and Fig. 6 using Eqs. (18) (19) (20).
@@
Figure 6: (color online) In the KR field, we present the relationship between the number of photon trajectories $n$ with different $l$-values and the collision parameter $b_{1}$ around the ATSW (with $Q=0.3$). The left figure shows the trajectory count for $n_{1}$, while the right figure illustrates the trajectory counts for $n_{2}$ (solid lines) and $n_{3}$ (dashed lines). We analyze the trajectory counts for photons under different Lorentz-violating parameters $l$: $l=0.01$ (black lines), $l=0.05$ (blue lines), and $l=0.1$ (red lines). The parameters are set as $R=2.6$, $M_{1}=1$, and $M_{2}=1.2$.
Fig. 5  and Fig. 6  illustrate the relationship between the number of photon trajectories and the impact parameter $b_{1}$ around an ATSW, under the influence of charge $Q$ and the Lorentz-violating parameter $l$. In Fig. 5 (a) and  6 (a), it is observed that the photon trajectories for $n_{1}$ are similar to those of a black hole, indicating that these photons will continue to propagate in spacetime $M_{1}$. In contrast, trajectories corresponding to $n_{2}$ and $n_{3}$ appear only in the ATSW. When a photon enters the throat from spacetime $M_{1}$, passes through the potential barrier in spacetime $M_{2}$, and returns to spacetime $M_{1}$, the outgoing photons intersect with the back side of the thin accretion disk if the conditions in Fig. 5 (b) and  6 (b) are met—specifically, when $n_{2}<0.75$ (solid line) and $0.75<n_{3}$ (dashed line). Conversely, when $n_{2}<1.25$ and $1.25<n_{3}$, the outgoing photons intersect with the front side of the thin accretion disk. Furthermore, analysis of Fig. 5 (b) and  6 (b) indicates that the range of the impact parameter $b_{1}$ decreases as both the charge $Q$ and Lorentz-violating parameter $l$ increase. This result suggests that additional photon rings contract toward the inner region of the ATSW shadow with increasing values of $Q$ and $l$. Additionally, it is noted that $n_{1}$, $n_{2}$, and $n_{3}$ are more significantly influenced by the Lorentz-violating parameter $l$ than by the charge $Q$.
@@ III.2 Observed intensity and transfer function
$\frac{I_{\nu^{\prime}}^{\mathrm{obs}}}{I_{\nu}^{\mathrm{em}}}=\frac{\nu^{\prime 3}}{\nu^{3}},$ (21)
here, $\nu^{\prime}$ refers to the redshifted frequency, denoted as $\nu^{\prime}=\sqrt{f}\nu$, which consequently yields the observed intensity
@@
Based on the Eq. (22), we integrated the observed intensity to obtain the total observed intensity at the intersection points
$I^{\mathrm{obs}}=\int I_{\nu^{\prime}}^{\mathrm{obs}}\mathrm{~d}\nu^{\prime}=\int f^{2}I_{\nu}^{\mathrm{em}}\mathrm{~d}v=f^{2}(r)I^{\mathrm{em}}(r),$ (23)
here, $I^{\mathrm{em}}(r)=\int I_{\nu}^{\mathrm{em}}\mathrm{d}\nu$ represents the total intensity radiated from the accretion disk. Consequently, we obtain the total observed intensity from all intersection points
$I_{\mathrm{obs}}(b)=\left.\sum_{m}I^{\mathrm{em}}(r)f^{2}(r)\right|_{r=r_{m}\left(b_{1}\right)},$ (24)
with $r_{m}\left(b_{1}\right)$ represents the transfer function, and $b_{1}$ denotes the impact parameter. The radial coordinate of the intersection point between the thin accretion disk and the photon trajectory is determined by the transfer function $r_{m}\left(b_{1}\right)$. When $m=1$, $m=2$, and $m=3$, they correspond to the direct emission, lensing ring, and photon ring of the thin accretion disk, respectively. The relationship between the transfer function and the impact parameter is illustrated in Fig. 7  and Fig. 8 , where we have considered the effects of varying charge and Lorentz -violating parameters.
@@
Fig. 7  and Fig. 8  show the relationship between the transfer function and the impact parameter $b_{1}$ under the influence of the charge $Q$ and the Lorentz-violating parameter $l$. The selected values for $Q$ are $0.1$, $0.3$, and $0.5$, while for $l$ they are $0.01$, $0.05$, and $0.1$. From Fig. 7  and Fig. 8 , it can be observed that the first transfer function ($m=1$), represented by the black solid line, corresponds to the direct image of the accretion disk. Due to its relatively small slope-where $\mathrm{d}r(b_{1})/\mathrm{d}b_{1}$ is defined as the demagnification factor, the small demagnification leads to the observed intensity being dominated by the direct image. The second transfer function ($m=2$), shown as the blue solid line, exhibits a larger demagnification factor, forming a demagnified lensing ring from the back side of the thin accretion disk and contributing only marginally to the observed intensity. The third transfer function ($m=3$), represented by the red solid line, displays the largest demagnification factor, producing a significantly demagnified photon ring from the front side of the thin accretion disk.
As shown in Fig. 7  and Fig. 8 , a comparison of the transfer functions between the BH and the ATSW reveals that the ATSW exhibits two additional transfer functions near the critical curves $Zb_{c_{2}}$ and $b_{c_{1}}$, represented by the blue and red dashed lines in the figures. The slopes of the new second and third transfer functions are approximately the same as those of the conventional second and third transfer functions, respectively. In summary, the additional transfer functions in the ATSW produce extra light rings, a feature that distinguishes it from the BH. Furthermore, we observe that as the charge $Q$ and the Lorentz-violating parameter $l$ increase, the positions of all transfer functions shift inward, with the second and third transfer functions undergoing more significant changes. It is also noted that the Lorentz-violating parameter $l$ has a more pronounced influence on the transfer functions than the charge $Q$.
@@
To better understand the optical appearance of an ATSW in a KR field, this section investigates its observable features through two distinct emission models of a thin accretion disk. The optical appearance of the black hole observed by a static observer at infinity is also determined by the emission model of the thin accretion diskWang:2025opr .
@@
$I_{A}^{\mathrm{em}}(r)=\begin{cases}0&r<r_{isco},\\ \left(\frac{1}{r-\left(r_{isco}-1\right)}\right)^{2}&r\geq r_{isco},\end{cases}$ (25)
here, $r_{isco}$ represents the photon unstable orbit in the thin accretion disk, defined as
@@
The radiation peak is achieved at $r=r_{isco}$, and then decreases as $r$ increases, approaching zero. There is no radiation emission within the photon unstable orbit radius. We have plotted the radiation function, as shown in Fig. 9 (a). Additionally, we have depicted the optical appearance of black holes with the same mass parameter and radiation model $A$ in the first row of Fig. 10. For better comparison, we have performed numerical simulations of the observational intensity, density distribution, and local density distribution of ATSW using model $A$, as shown in the second row of Fig. 10 . By comparing Fig. 10 (a) and Fig. 10 (b), it is evident that BH and ATSW exhibit significant differences in their optical structures, with distinct spatial separations between the direct radiation ring, lens ring, and photon ring. Observing Fig. 10 (a), we find that the observational intensity of ATSW appears at the critical curve $b_{1}\simeq 6.8289M_{1}$, with an intensity of $0.4529$, which subsequently decreases to zero. The lens ring is located within a narrow region between the critical curves $b_{1}\simeq 5.3774M_{1}$ and $b_{1}\simeq 5.8891M_{1}$. The photon rings appear near the critical curves at $b_{1}\simeq 3.7221M_{1}$, $b_{1}\simeq 4.9676M_{1}$, and $b_{1}\simeq 5.1213M_{1}$. Compared to BH, intriguingly, two new photon rings emerge in the shadow of ATSW, located near the critical curves at $b_{1}\simeq 3.7221M_{1}$ and $b_{1}\simeq 4.9676M_{1}$, as shown in Fig. 10 (a) and  10 (d). By examining Fig. 10 (b) and  10 (c), it is observed that the direct emission ring is near the edge of the black disk, while the lens ring is positioned inside the black disk. Furthermore, Fig. 10 (e) and  10 (f) reveal that the additional photon rings of ATSW are closer to the black disk.
@@
$I_{2}^{\mathrm{em}}(r)=\begin{cases}0&r<r_{h},\\ \frac{\frac{\pi}{2}-\tan^{-1}\left(r-\left(r_{\mathrm{isco}}-1\right)\right)}{\frac{\pi}{2}-\tan^{-1}\left(r_{ph}\right)}&r\geq r_{h},\end{cases}$ (27)
in which, $r_{h}$ represents the event horizon radius. The radiation peak is achieved at $r=r_{h}$, and then decreases as $r$ increases, approaching zero. There is no radiation emission within the photon unstable orbit. The image of radiation model $B$ is shown in Fig. 9 (b), where the decline in radiation value with increasing r is more gradual compared to radiation model $A$. Additionally, we have plotted the optical appearance of black holes with the same mass parameter and radiation model B in the first row of Fig. 11 . For better comparison, we have performed numerical simulations of the observational intensity, density distribution, and local density distribution of ATSW using model $B$, as shown in the second row of Figure 10. It can be observed that there is partial overlap in the regions of the direct radiation ring, lens ring, and photon ring in ATSW. The direct ring in ATSW starts radiating at $b_{1}\simeq 2.7449M_{1}$. Interestingly, the lens ring encompasses the photon ring, collectively forming a multi-layered bright ring feature, as illustrated in Fig. 11 (e) and  (f).
@@
We have plotted the observational intensity, density maps, and local density maps for the ATSW in the KR field under two sets of parameters: $Q=0.3$, $l=0.01$ and $Q=0.3$, $l=0.05$, for two radiation models, as shown in Fig. 12 .Observing Fig. 10 . 11  and  12 , it is found that in radiation model $A$, for the case of $Q=0.3$, $l=0.01$, the two additional photon rings are located at $b_{1}\simeq 3.8425M_{1}$ and $b_{1}\simeq 4.0561M_{1}$, respectively; for $Q=0.3$, $l=0.05$, the two additional photon rings are located at $b_{1}\simeq.9942M_{1}$ and $b_{1}\simeq 4.7456M_{1}$, respectively. In radiation model $B$, for $Q=0.3$, $l=0.01$, the additional lensing ring is located between $b_{1}\simeq 3.8221M_{1}$ and $b_{1}\simeq 5.056M_{1}$; for $Q=0.3$, $l=0.05$, the additional lensing ring is located between $b_{1}\simeq 3.9686M_{1}$ and $b_{1}\simeq 5.0764M_{1}$. The results indicate that as the charge $Q$ and the Lorentz-violating parameter $l$ increase, the range of light bands outside the shadow decreases, but the range of specific additional rings increases. Therefore, the trend of change in the external shadow region is the same as that of the KR field BH, while the trend of change in specific additional rings is opposite.

### v3

@@ II Null Geodesics and Effective Potential of the ATSW in the Kalb-Ramond Field
With Eqs. (19) (20) (21), we can plot the photon trajectories for different values of the charge $Q$ and the Lorentz-violation parameter $l$, as shown in Figs. 5  and  6 . Taking $Q=0.1$ and $l=0.01$ as an example, Fig.  5 (a), (b), and (c) illustrate three typical types of photon motion. For the case where $Zb_{c_{2}}<b_{1}<b_{c_{1}}$, photons originating from infinity in spacetime $M_{1}$ exhibit an increasing number of trajectory loops in spacetime $M_{2}$ as the impact parameter $b_{i}$ decreases. By comparing the corresponding columns in Figs. 5  and  6 , we observe that $b_{c_{1}}$ decreases with an increase in either $Q$ or $l$, while $Zb_{c_{2}}$ shows the opposite trend. Furthermore, it is evident that the photon trajectories are more significantly influenced by the Lorentz-violation parameter $l$.
@@ III.1 The trajectory of a photon
First, in a stationary coordinate system with the BH fixed in the equatorial plane, consider a static observer at infinite distance in the direction of the North Pole. We assume that the light source surrounding the black hole is a geometrically and optically thin accretion disk situated on the equatorial plane. Based on the backward ray-tracing approach, if the light rays emitted by the observer intersect the accretion disk, the light emanating from the corresponding position on the disk can be detected by the observer. The total number of orbits, as a function of $b$, which can be expressed as
@@
Here, $n_{2}$ and $n_{3}$ serve as additional orbit functions for the ATSW, corresponding to additional photons. This situation is analogous to that of a BH. To better illustrate the relationship between the number of photon trajectories and the impact parameter, we plotted Figs. 7  and  8 using Eqs. (26) (27) and (28).
@@
Figure 8: (color online) In the KR field, we present the relationship between the number of photon trajectories $n$ with different $l$-values and the collision parameter $b_{1}$ around the ATSW (with $Q=0.3$). The left figure shows the trajectory count for $n_{1}$, while the right figure illustrates the trajectory counts for $n_{2}$ (solid lines) and $n_{3}$ (dashed lines). We analyze the trajectory counts for photons under different Lorentz-violation parameters $l$: $l=0.01$ (black lines), $l=0.05$ (blue lines), and $l=0.1$ (red lines). The parameters are set as $R=2.6$, $M_{1}=1$, and $M_{2}=1.2$.
Figs. 7  and  8  illustrate the relationship between the number of photon trajectories and the impact parameter $b_{1}$ around an ATSW, under the influence of charge $Q$ and the Lorentz-violation parameter $l$. In Figs. 7 (a) and  8 (a), it is observed that the photon trajectories for $n_{1}$ are similar to those of a black hole, indicating that these photons will continue to propagate in spacetime $M_{1}$. In contrast, trajectories corresponding to $n_{2}$ and $n_{3}$ appear only in the ATSW. When a photon enters the throat from spacetime $M_{1}$, passes through the potential barrier in spacetime $M_{2}$, and returns to spacetime $M_{1}$, the outgoing photons intersect with the back side of the thin accretion disk if the conditions in Figs. 7 (b) and  8 (b) are met—specifically, when $n_{2}<0.75$ (solid line) and $0.75<n_{3}$ (dashed line). Conversely, when $n_{2}<1.25$ and $1.25<n_{3}$, the outgoing photons intersect with the front side of the thin accretion disk. Furthermore, analysis of Figs. 7 (b) and  8 (b) indicates that the range of the impact parameter $b_{1}$ decreases as both the charge $Q$ and Lorentz-violation parameter $l$ increase. This result suggests that additional photon rings contract toward the inner region of the ATSW shadow with increasing values of $Q$ and $l$. Additionally, it is noted that $n_{1}$, $n_{2}$, and $n_{3}$ are more significantly influenced by the Lorentz-violation parameter $l$ than by the charge $Q$.
@@ III.2 Observed intensity and transfer function
$\frac{I_{\nu^{\prime}}^{\mathrm{obs}}}{I_{\nu}^{\mathrm{em}}}=\frac{\nu^{\prime 3}}{\nu^{3}}.$ (29)
Here, $\nu^{\prime}$ refers to the redshifted frequency. Under the assumption of a static emitter, this is denoted as $\nu^{\prime}=\sqrt{f}\nu$, which consequently yields the observed intensity
@@
The frequency shift factor thus accounts solely for gravitational redshift in this context. Based on the Eq. (30), we integrated the observed intensity to obtain the total observed intensity at the intersection points
$I^{\mathrm{obs}}=\int I_{\nu^{\prime}}^{\mathrm{obs}}\mathrm{~d}\nu^{\prime}=\int f^{2}I_{\nu}^{\mathrm{em}}\mathrm{~d}v=f^{2}(r)I^{\mathrm{em}}(r).$ (31)
Here, $I^{\mathrm{em}}(r)=\int I_{\nu}^{\mathrm{em}}\mathrm{d}\nu$ represents the total intensity radiated from the accretion disk. Consequently, we obtain the total observed intensity from all intersection points as a function of the impact parameter $b_{1}$
$I_{\mathrm{obs}}(b_{1})=\left.\sum_{m}I^{\mathrm{em}}(r)f^{2}(r)\right|_{r=r_{m}\left(b_{1}\right)}.$ (32)
With $r_{m}\left(b_{1}\right)$ represents the transfer function, and $b_{1}$ denotes the impact parameter. The radial coordinate of the intersection point between the thin accretion disk and the photon trajectory is determined by the transfer function $r_{m}\left(b_{1}\right)$. When $m=1$, $m=2$, and $m=3$, they correspond to the direct emission, lensing ring, and photon ring of the thin accretion disk, respectively. The relationship between the transfer function and the impact parameter is illustrated in Figs. 9  and 10 , where we have considered the effects of varying charge and Lorentz -violation parameters.
@@
Figs. 9  and  10  show the relationship between the transfer function and the impact parameter $b_{1}$ under the influence of the charge $Q$ and the Lorentz-violation parameter $l$. The selected values for $Q$ are $0.1$, $0.3$, and $0.5$, while for $l$ they are $0.01$, $0.05$, and $0.1$. From Figs. 9  and  10 , it can be observed that the first transfer function ($m=1$), represented by the black solid line, corresponds to the direct image of the accretion disk. Due to its relatively small slope-where $\mathrm{d}r(b_{1})/\mathrm{d}b_{1}$ is defined as the demagnification factor, the small demagnification leads to the observed intensity being dominated by the direct image. The second transfer function ($m=2$), shown as the blue solid line, exhibits a larger demagnification factor, forming a demagnified lensing ring from the back side of the thin accretion disk and contributing only marginally to the observed intensity. The third transfer function ($m=3$), represented by the red solid line, displays the largest demagnification factor, producing a significantly demagnified photon ring from the front side of the thin accretion disk.
As shown in Figs. 9  and  10 , a comparison of the transfer functions between the BH and the ATSW reveals that the ATSW exhibits two additional transfer functions near the critical curves $Zb_{c_{2}}$ and $b_{c_{1}}$, represented by the blue and red dashed lines in the figures. The slopes of the new second and third transfer functions are approximately the same as those of the conventional second and third transfer functions, respectively. In summary, the additional transfer functions in the ATSW produce extra light rings, a feature that distinguishes it from the BH. Furthermore, we observe that as the charge $Q$ and the Lorentz-violation parameter $l$ increase, the positions of all transfer functions shift inward, with the second and third transfer functions undergoing more significant changes. It is also noted that the Lorentz-violation parameter $l$ has a more pronounced influence on the transfer functions than the charge $Q$.
@@
To better understand the optical appearance of the ATSW in a KR field, this section investigates its observable features through two distinct emission models of a thin accretion disk. The optical appearance of the black hole observed by a static observer at infinity is also determined by the emission model of the thin accretion diskWang:2025opr .
@@
$I_{A}^{\mathrm{em}}(r)=\begin{cases}0&r<r_{isco},\\ \left(\frac{1}{r-\left(r_{isco}-1\right)}\right)^{2}&r\geq r_{isco}.\end{cases}$ (33)
Here, $r_{isco}$ denotes the innermost stable circular orbit (ISCO) of massive particles, defined as
@@
The radiation flux peaks at $r=r_{isco}$ and then decreases with radius, asymptotically approaching zero. For an observer at infinity, the effective radiation originating from within the photon unstable orbit is negligible, as any emission from this region is either highly redshifted or captured by the central object. We have plotted the radiation function, as shown in Fig. 11 (a). Additionally, we have depicted the optical appearance of black holes with the same mass parameter and radiation model $A$ in the first row of Fig. 12 . For better comparison, we have computed the observed intensity and constructed the corresponding intensity distribution and local intensity maps for the ATSW using Model A through our ray-tracing method. The results are shown in the second row of Fig. 12 . By comparing Figs. 12 (a) and  12 (b), it is evident that BH and ATSW exhibit significant differences in their optical structures, with distinct spatial separations between the direct radiation ring, lens ring, and photon ring. Observing Fig. 12 (a), we find that the observational intensity of ATSW appears at the critical curve $b_{1}\simeq 6.8289M_{1}$, with an intensity of $0.4529$, which subsequently decreases to zero. The lens ring is located within a narrow region between the critical curves $b_{1}\simeq 5.3774M_{1}$ and $b_{1}\simeq 5.8891M_{1}$. The photon rings appear near the critical curves at $b_{1}\simeq 3.7221M_{1}$, $b_{1}\simeq 4.9676M_{1}$, and $b_{1}\simeq 5.1213M_{1}$. Compared to BH, intriguingly, two new photon rings emerge in the shadow of ATSW, located near the critical curves at $b_{1}\simeq 3.7221M_{1}$ and $b_{1}\simeq 4.9676M_{1}$, as shown in Figs. 12 (a) and  12 (d). By examining Figs. 12 (b) and  12 (c), we observe that the direct emission ring is near the edge of the black disk, while the lens ring is positioned inside the black disk. Furthermore, Figs. 12 (e) and  12 (f) reveal that the additional photon rings of ATSW are closer to the black disk.
@@
$I_{2}^{\mathrm{em}}(r)=\begin{cases}0&r<r_{h},\\ \frac{\frac{\pi}{2}-\tan^{-1}\left(r-\left(r_{\mathrm{isco}}-1\right)\right)}{\frac{\pi}{2}-\tan^{-1}\left(r_{ph}\right)}&r\geq r_{h}.\end{cases}$ (35)
In which, $r_{h}$ represents the event horizon radius. The radiation peak is achieved at $r=r_{h}$, and then decreases as $r$ increases, approaching zero. The image of radiation model $B$ is shown in Fig. 11 (b), where the decline in radiation value with increasing r is more gradual compared to radiation model $A$. Additionally, we have plotted the optical appearance of black holes with the same mass parameter and radiation model B in the first row of Fig. 13 . For better comparison, we have adopted a numerical method to calculate the observational intensity, intensity distribution, and local intensity distribution of an ATSW using model $B$, as shown in the second row of Figure 10. It can be observed that there is partial overlap in the regions of the direct radiation ring, lens ring, and photon ring in the ATSW. The direct ring in the ATSW starts radiating at $b_{1}\simeq 2.7449M_{1}$. Interestingly, the lens ring encompasses the photon ring, collectively forming a multi-layered bright ring feature, as illustrated in Figs. 13 (e) and  13 (f).
@@
We have plotted the observational intensity, density maps, and local density maps for the ATSW in the KR field under two sets of parameters: $Q=0.3$, $l=0.01$ and $Q=0.3$, $l=0.05$, for two radiation models, as shown in Fig. 14 .Observing Figs. 12 . 13  and  14 , we find that in radiation model $A$, for the case of $Q=0.3$, $l=0.01$, the two additional photon rings are located at $b_{1}\simeq 3.8425M_{1}$ and $b_{1}\simeq 4.0561M_{1}$, respectively; for $Q=0.3$, $l=0.05$, the two additional photon rings are located at $b_{1}\simeq.9942M_{1}$ and $b_{1}\simeq 4.7456M_{1}$, respectively. In radiation model $B$, for $Q=0.3$, $l=0.01$, the additional lensing ring is located between $b_{1}\simeq 3.8221M_{1}$ and $b_{1}\simeq 5.056M_{1}$; for $Q=0.3$, $l=0.05$, the additional lensing ring is located between $b_{1}\simeq 3.9686M_{1}$ and $b_{1}\simeq 5.0764M_{1}$. The results indicate that as the charge $Q$ and the Lorentz-violation parameter $l$ increase, the range of light bands outside the shadow decreases, but the range of specific additional rings increases. Therefore, the trend of change in the external shadow region is the same as that of the KR field BH, while the trend of change in specific additional rings is opposite.